# Rydberg superatoms: An artificial quantum system for quantum information processing and quantum optics

### v1

@@ Abstract
Dense Rydberg atom ensembles display intriguing collective behaviors mediated by their strong, long-range dipole-dipole interactions. These collective effects, often modeled using Rydberg superatoms, have gained significant attention across various fields due to their potential applications in quantum information processing and quantum optics. In this review article, we delve into the theoretical foundations of Rydberg interactions and explore experimental techniques for their manipulation and detection. We also discuss the latest advancements in harnessing Rydberg collective effects for quantum computation and optical quantum technologies. By synthesizing insights from theoretical studies and experimental demonstrations, we aim to provide a comprehensive overview of this rapidly evolving field and its potential impact on the future of quantum technologies.
@@ I INTRODUCTION
Rydberg atoms exhibit striking properties that arise from their exaggerated atomic size and strong interatomic interactions.Gallagher (1988, 1994); Saffman, Walker, and Mølmer (2010) The large electric dipole moment couples with external electric fields, allowing the manipulation of the electronic as well as motional states of Rydberg atoms with laser and microwave (MW) fields. On the other hand, strong and long-range dipole-dipole or van der Waals interactions between Rydberg atoms are found even when they are separated by macroscopic distances.Urban et al. (2009); Gaëtan et al. (2009) Such interactions, moreover, can be controlled additionally by laser or MW fields. This paves the way for realizing scalable quantum information processing architectures, Jaksch et al. (2000) and exploring novel many-body states Browaeys and Lahaye (2020) and nonlinear quantum optics.Firstenberg, Adams, and Hofferberth (2016)
@@
At the heart of the Rydberg physics lies strong interaction-induced collective behaviors, Heidemann et al. (2007); Dudin et al. (2012); Zeiher et al. (2015) such as the Rydberg blockade.Lukin et al. (2001) In high-lying Rydberg states, the interatomic interaction ($1\sim 100$ MHz) can be much stronger than typical laser Rabi frequencies (tens of kHz $\sim$ a few MHz). As a result, the laser excitation of one atom inhibits the excitation of neighboring atoms within a certain distance (i.e. blockade radius). Rydberg blockade becomes the foundation of many quantum gate schemes.Jaksch et al. (2000); Lukin et al. (2001); Saffman, Walker, and Mølmer (2010) Within the blockade radius, the many-body ground state and singly excited Rydberg state can be excited. The former consists of atoms in the electronic ground state, while the latter is a collective state where one atom is in the Rydberg state while all other atoms are in the electronic ground state. Dynamics of such collective states are conveniently described by an artificial superatom.Lukin et al. (2001) This description can be extended to many-atom settings, where the collective effect allows for the creation of exotic many-body correlated states.Pohl, Demler, and Lukin (2010); Olmos, González-Férez, and Lesanovsky (2009); Honer et al. (2010); Weimer et al. (2010); Ji, Ates, and Lesanovsky (2011); de Léséleuc et al. (2019) These potentials make Rydberg atoms a leading candidate for implementing quantum simulation and quantum computing tasks. Moreover, strong Rydberg atom interactions can be mapped to optical fields, generating sizable optical nonlinearities.Pritchard et al. (2010) These have found wide usage in exploring nonlinear quantum optics Peyronel et al. (2012); Firstenberg, Adams, and Hofferberth (2016) and realizing optical quantum devices, such as single-photon sources,Dudin and Kuzmich (2012); Ripka et al. (2018); Padrón-Brito et al. (2021); Ornelas-Huerta et al. (2020); Shi et al. (2022) photonic quantum logic gates,Tiarks et al. (2016, 2019); Stolz et al. (2022) photonic entanglement filter,Ye et al. (2023) single-photon absorber,Tresp et al. (2016) single-photon switch,Baur et al. (2014) and transistors.Baur et al. (2014); Gorniaczyk et al. (2014)
@@
In this review article, we will delve into the theory underpinning of the Rydberg superatom, related collective effects and key experimental achievements. In particular, we will review the harness of Rydberg collectivity in two critical areas: quantum information processing and quantum optics. We start with the introduction of Rydberg atoms and Rydberg superatoms in Sec. II. The atomic properties and two-body interactions of Rydberg atoms, and the emergence of the collective states described by Rydberg superatoms, are discussed. In Sec. III, coherent laser excitation of Rydberg atoms and the preparation of Rydberg superatoms are given. Quantum interference of Rydberg superatoms is discussed. In Sec. IV, quantum information processing based on Rydberg superatoms and collective effects are discussed. We focus on the creation of quantum logic gates, entanglements, and simulations using Rydberg atoms. We describe the application of the Rydberg atom systems in quantum optics in Sec. V, including the demonstration and realization of strong nonlocal nonlinear optics, single-photon applications, and hybridizing quantum systems with Rydberg interactions and superatoms. Finally, we conclude in Sec. VI.
@@ II.1 Characteristics of Rydberg atoms
${}^{85}$b $nS_{1/2}$ 3.1311804Li et al. (2003) 0.1784Li et al. (2003) -1.8Lorenzen and Niemax (1983)
${}^{85}$b $nP_{3/2}$ 2.64142Li et al. (2021) 0.295Li et al. (2021) 0.97495Lorenzen and Niemax (1983) 14.6001Lorenzen and Niemax (1983)
${}^{85}$b $nD_{5/2}$ 1.3464572Li et al. (2003) -0.596Li et al. (2003) -1.50517Lorenzen and Niemax (1983) -2.4206Lorenzen and Niemax (1983)
${}^{85}$b $nF_{7/2}$ 0.016411Li et al. (2021) -0.0784Li et al. (2021) -0.36005Lorenzen and Niemax (1983) 3.2390Lorenzen and Niemax (1983)
${}^{133}$s $nS_{1/2}$ 4.0493532Deiglmayr et al. (2016) 0.23915Deiglmayr et al. (2016) 0.06Deiglmayr et al. (2016) 11Deiglmayr et al. (2016)
${}^{133}$s $nP_{3/2}$ 3.5590676Deiglmayr et al. (2016) 0.37469Deiglmayr et al. (2016) -0.67431Lorenzen and Niemax (1984) 22.3531Lorenzen and Niemax (1984)
${}^{133}$s $nD_{5/2}$ 2.4663144Deiglmayr et al. (2016) 0.01381Deiglmayr et al. (2016) -0.392Deiglmayr et al. (2016) -1.9Deiglmayr et al. (2016)
${}^{133}$s $nF_{7/2}$ 0.0335646Bai et al. (2023) -0.2052Bai et al. (2023)
Starting from the expression for Rydberg’s binding energy, we can obtain a series of scaling laws that exhibit exaggerated characteristics concerning the principal quantum number $n$, or more precisely, the effective principal quantum number $n^{*}=(n-\delta_{n,l,j})$. The energy spacing between adjacent Rydberg levels is given by
@@
which follows a $n^{*-3}$ scaling. The corresponding transition wavelength typically falls within the millimeter-wave band between 10 mm (30 GHz) and 1 mm (300 GHz). This feature offers exceptional precision in spectroscopic measurements, Prozument et al. (2011); Gordon et al. (2014) showing transient amplifiers of microwave radiation Raimond et al. (1982); Moi et al. (1983); Goy et al. (1983) and contributing to the development of the fifth generation mobile communication network.Deb and Kjærgaard (2018); Yuan et al. (2023a) It serves a crucial role in the advancement of atomic physics and quantum science research.Carter and Martin (2013); Kiffner et al. (2016); Grimes et al. (2017); Wade et al. (2017); Peper et al. (2019); Suleymanzade et al. (2020); Kanungo et al. (2022); Legaie, Raithel, and Anderson (2023)
Figure 1: (a) The numerical data demonstrates the scaling laws for the average orbit radius of Rydberg atoms in the $S$ orbit (blue) and circular orbit (red), expressed in terms of the effective principal quantum number denoted as $n^{*}$ for each case. (b) The scaling laws govern the radial dipole matrix elements for transitions between neighboring Rydberg states, denoted as $|\langle nP_{3/2}|e\hat{r}|nS_{1/2}\rangle|$ (blue), as well as for transitions between a low-lying state and a Rydberg state, represented as $|\langle nP_{3/2}|e\hat{r}|5S_{1/2}\rangle|$ (red), where $n^{*}=n-\delta_{n,2,j}$ for the $P$ orbit.
The average radius of the electron orbit, determined by the radial wave function for a Rydberg state, is given by,
@@
where $a_{0}\approx 5.29\times 10^{-2}~{}\rm nm$ represents the Bohr radius.Gallagher (1994); Theodosiou (1984) This formula offers an exact description of the radius of Rydberg atomic orbitals with high principal quantum numbers, as the Coulomb potential predominates in this regime. The Rydberg atomic-wave function, influenced by quantum defects, deviates from that of the hydrogen atom mainly by a phase shift. As a result, the average value of physical quantities can be directly derived from the hydrogen atom formula, with the only alteration being the substitution of the principal quantum number $n$ with the effective principal quantum number $n^{*}$.Kazimierczuk et al. (2014); Heckötter et al. (2017) For low $l$ states, particularly $l=0$ ($S$ states), the average radius simplifies to be $23n^{*2}a_{0}/$ 2, but for the circular orbit with $l=n-1$, where the quantum defect is zero, the value becomes $(n^{2}+n/2)a_{0}$. For large $n$ the radius of the Rydberg electron can be hundreds of nanometers. This permits the formation of ultra-long range Rydberg molecules, Herzberg (1987); Greene, Dickinson, and Sadeghpour (2000); Bendkowsky et al. (2009, 2010); Kiffner et al. (2012); Krupp et al. (2014); Camargo et al. (2016); Li et al. ; Eiles and Greene (2017); Eiles et al. (2017); Shaffer, Rittenhouse, and Sadeghpour (2018); Whalen et al. (2020); Giannakeas et al. (2020); Peper and Deiglmayr (2021); Christian Fey and Schmelcher (2020); Hummel, Eiles, and Schmelcher (2021) where ground state atoms are trapped within the Rydberg electron cloud. In Fig. 1(a), the ARC toolbox Šibalić et al. (2017) is used to perform numerical calculation of the average radii for the $S$ and circular orbits of Rubidium, encompassing the principal quantum number $n$ from 10 to 100. The data exhibit quadratic scaling with $n^{*}$, closely aligning with the analytical expression.
@@
The spontaneous decay rate from $|n,l,j\rangle$ to lower lying state $|n^{\prime},l^{\prime},j^{\prime}\rangle$ considering the fine structure in SI units is governed by the Einstein coefficient:
@@
where $\omega_{n^{\prime}l^{\prime}j^{\prime},nlj}$ designates the corresponding transition angular frequency, $l_{max}$ signifies the larger of the two values between $l$ and $l^{\prime}$, and the curly braces indicate a Wigner-$6j$ symbol.Seiler et al. (2016); Beterov et al. (2009) For low angular momentum states, Rydberg states can decay to many low-lying states through dipole transitions. At zero temperature, the decay rate is proportional to $n^{*-3}$. However, in circular Rydberg states ($l=n-1$), they can only undergo a transition to an adjacent Rydberg state via a dipole-allowed process.Hulet and Kleppner (1983); Nguyen et al. (2018) This leads to the spontaneous emission rate of the circular state being proportional to $n^{*-5}$. In all the situations, long lifetimes of Rydberg atoms are found, which are important for maintaining quantum coherence, a crucial factor for quantum computation and simulation.
@@
Compared to the high $l$ states, the quantum defects of the low $l\leq 3$ states have a greater impact on the energy levels correction. Consequently, the weak external electric field will not cause the Rydberg atom in the low orbit state to endure a direct transition between the adjacent orbits but will instead present the energy level shift as a second-order perturbation, which is in the form of a quadratic Stark effectKhadjavi, Lurio, and Happer (1968)
@@
$\Delta E_{\rm Stark}=\sum_{n^{\prime},l^{\prime}j^{\prime}\neq n,l,j}\frac{|\langle n^{\prime},l^{\prime},j^{\prime},m_{j}^{\prime}|e\hat{r}\cdot\bm{E}|n,l,j,m_{j}\rangle|^{2}}{E_{n^{\prime},l^{\prime},j^{\prime}}-E_{n,l,j}}=\frac{1}{2}\alpha_{0}E^{2}$ (9)
@@
where we have assumed a static electric field $E$ is applied along the $z$-axis creating off-diagonal couplings between states, with the selection rule $\Delta l=\pm 1$, $\Delta m_{j}=0$. As determined by the Wigner-Eckart theorem,Weissbluth (1978) the dipole matrix elements are linked to the radial matrix elements $\langle n^{\prime},l^{\prime},j^{\prime}|e\hat{r}|{n},l,j\rangle$ through coefficients independent of the principal quantum number. Therefore, the scalar polarizability $\alpha_{0}$ scales with $n^{*7}$ due to the combined scaling of the radial dipole matrix element and the energy spacing for neighboring Rydberg states in the denominator.Zimmerman et al. (1979); O’Sullivan and Stoicheff (1985); Lai et al. (2018); Bai et al. (2020a); Chakraborty and Sahoo (2023) It should be noted that states possessing quantum numbers of angular momentum $l>3$ manifest degeneracy, leading to a linear Stark shift. For a more precise description of the Stark structure of the Rydberg states in alkali atoms, direct diagonalization of the energy matrix is required. The high polarizability of Rydberg atoms can be exploited as highly accurate sensors for electric fieldsSedlacek et al. (2012); Fan et al. (2014, 2015); Anderson, Paradis, and Raithel (2018); Cox et al. (2018); Meyer et al. (2020); Jing et al. (2020); Crawford et al. (2021); Meyer et al. (2021); Calero et al. (2021); Hu et al. (2022); Liu et al. (2022); Fancher et al. (2023); Liu et al. (2023a); Yuan et al. (2023b) and the nondestructive imaging of atoms and ions.Günter et al. (2012); Gross, Vogt, and Li (2020)
@@
Table 2: Scaling of various properties of Rydberg atoms with the effective principal quantum numbers $n^{*}$. The data in the parentheses are scaling in the circular state.

### v2

@@ Abstract
Dense atom ensembles with Rydberg excitations display intriguing collective effects mediated by their strong, long-range dipole-dipole interactions. These collective effects, often modeled using Rydberg superatoms, have gained significant attention across various fields due to their potential applications in quantum information processing and quantum optics. In this review article, we delve into the theoretical foundations of Rydberg interactions and explore experimental techniques for their manipulation and detection. We also discuss the latest advancements in harnessing Rydberg collective effects for quantum computation and optical quantum technologies. By synthesizing insights from theoretical studies and experimental demonstrations, we aim to provide a comprehensive overview of this rapidly evolving field and its potential impact on the future of quantum technologies.
@@ I INTRODUCTION
Rydberg atoms exhibit striking properties that arise from their large atomic size and strong interatomic interactions.Gallagher (1988, 1994); Saffman, Walker, and Mølmer (2010) The huge induced electric dipole moment couples with external electric fields, allowing the manipulation of the electronic as well as motional states of Rydberg atoms with laser and microwave (MW) fields. Importantly, strong and long-range dipole-dipole or van der Waals (vdW) interactions are found between Rydberg atoms separated by macroscopic distances of over 10 micrometers.Urban et al. (2009); Gaëtan et al. (2009) Such interactions can be controlled by laser or MW fields. This paves the way for realizing scalable quantum information processing architectures, Jaksch et al. (2000) creating novel many-body states Browaeys and Lahaye (2020) and exploring nonlinear quantum optics.Firstenberg, Adams, and Hofferberth (2016)
@@
At the heart of the Rydberg physics lies strong interaction-induced collective behaviors.Heidemann et al. (2007); Dudin et al. (2012); Zeiher et al. (2015) In high-lying Rydberg states, the interatomic interaction (1111100100100100 MHz) can be much stronger than typical laser Rabi frequencies (tens of kHz–a few MHz).Saffman, Walker, and Mølmer (2010) As a result, the laser excitation of one atom inhibits the excitation of neighboring atoms within a certain distance (i.e. blockade radius). Within the blockade radius, only the many-body ground state and singly excited Rydberg state can be occupied. The former consists of atoms in the electronic ground state, while the latter is a collective state where one atom is in the Rydberg state while all other atoms are in the electronic ground state. Dynamics of such collective states are conveniently described by an artificial superatom.Lukin et al. (2001) This description can be extended to many-atom settings, where the collective effect allows for the creation of exotic many-body correlated states.Pohl, Demler, and Lukin (2010); Olmos, González-Férez, and Lesanovsky (2009); Honer et al. (2010); Weimer et al. (2010); Ji, Ates, and Lesanovsky (2011); de Léséleuc et al. (2019) These potentials make Rydberg atoms a leading candidate for implementing quantum simulation and quantum computing tasks. Moreover, strong Rydberg atom interactions can be mapped to optical fields, generating sizable optical nonlinearities.Pritchard et al. (2010) These have found wide usage in exploring nonlinear quantum optics Peyronel et al. (2012); Firstenberg, Adams, and Hofferberth (2016) and realizing optical quantum devices, such as single-photon sources,Dudin and Kuzmich (2012); Ripka et al. (2018); Padrón-Brito et al. (2021); Ornelas-Huerta et al. (2020); Shi et al. (2022) photonic quantum logic gates,Tiarks et al. (2016, 2019); Stolz et al. (2022) photonic entanglement filter,Ye et al. (2023) single-photon absorber,Tresp et al. (2016) single-photon switch,Baur et al. (2014) and single-photon transistors.Gorniaczyk et al. (2014)
In recent years, quite a few review papers have meticulously surveyed the rapidly evolving field of Rydberg atom technologies,Saffman, Walker, and Mølmer (2010); Browaeys and Lahaye (2020); Saffman (2016); Schauss (2018); Adams, Pritchard, and Shaffer (2020); Beterov et al. (2020); Henriet et al. (2020); Morgado and Whitlock (2021); Wu et al. (2021); Shi (2022); Yuan et al. (2023a) emphasizing their diverse applications and groundbreaking advancements in quantum computation, quantum simulation, and quantum sensing, showcasing the versatility and potential of Rydberg atoms. This review article distinguishes itself by concentrating specifically on the concept of Rydberg superatoms, with a particular emphasis on their role as a quantum interface between atoms and photons. Rydberg superatoms are created through collective excitation states, where multiple atoms collectively display characteristics akin to a single two-level atom. This collective behavior not only enhances the coherence and controllability of quantum states but also endows the interface with novel electromagnetic properties, enabling the creation and manipulation of photon-photon interactions at the quantum level. This review thus will contribute to the existing body of knowledge on Rydberg research, offering valuable insights and guidance for future studies.
@@
In this review article, we delve into the theory underpinning the Rydberg superatom, related collective effects, and key experimental achievements. In particular, we review the harnessing of Rydberg collectivity in two critical areas: quantum information processing and quantum optics. We start with an introduction to Rydberg atoms and Rydberg superatoms in Sec. II, discussing the atomic properties, two-body interactions of Rydberg atoms, and the emergence of collective states described by Rydberg superatoms. In Sec. III, coherent laser excitation of Rydberg atoms and the preparation of Rydberg superatoms are covered, including quantum interference of Rydberg superatoms. In Sec. IV, we discuss quantum information processing based on Rydberg superatoms, focusing on implementing quantum logic gates, generating quantum entanglement, and performing quantum simulations. We describe the application of Rydberg superatoms in quantum optics in Sec. V, including the demonstration and realization of strong nonlocal nonlinear optics, single-photon applications, and hybridizing quantum systems with Rydberg superatoms. Finally, we conclude in Sec. VI.
@@ II.1 Characteristics of Rydberg atoms
85Rb $nS_{1/2}$ 3.1311804Li et al. (2003) 0.1784Li et al. (2003) -1.8Lorenzen and Niemax (1983)
85Rb $nP_{3/2}$ 2.64142Li et al. (2021) 0.295Li et al. (2021) 0.97495Lorenzen and Niemax (1983) 14.6001Lorenzen and Niemax (1983)
85Rb $nD_{5/2}$ 1.3464572Li et al. (2003) -0.596Li et al. (2003) -1.50517Lorenzen and Niemax (1983) -2.4206Lorenzen and Niemax (1983)
85Rb $nF_{7/2}$ 0.016411Li et al. (2021) -0.0784Li et al. (2021) -0.36005Lorenzen and Niemax (1983) 3.2390Lorenzen and Niemax (1983)
133Cs $nS_{1/2}$ 4.0493532Deiglmayr et al. (2016) 0.23915Deiglmayr et al. (2016) 0.06Deiglmayr et al. (2016) 11Deiglmayr et al. (2016)
133Cs $nP_{3/2}$ 3.5590676Deiglmayr et al. (2016) 0.37469Deiglmayr et al. (2016) -0.67431Lorenzen and Niemax (1984) 22.3531Lorenzen and Niemax (1984)
133Cs $nD_{5/2}$ 2.4663144Deiglmayr et al. (2016) 0.01381Deiglmayr et al. (2016) -0.392Deiglmayr et al. (2016) -1.9Deiglmayr et al. (2016)
133Cs $nF_{7/2}$ 0.0335646Bai et al. (2023) -0.2052Bai et al. (2023)
Starting from the expression for Rydberg’s binding energy, we can obtain a series of scaling laws that exhibit outstanding characteristics concerning the principal quantum number $n$, or more precisely, the effective principal quantum number $n^{*}=(n-\delta_{n,l,j})$. The energy spacing between adjacent Rydberg levels is given by
@@
which follows a $n^{*-3}$ scaling. The corresponding transition wavelength typically falls within the millimeter-wave band between 10 mm (30 GHz) and 1 mm (300 GHz).Šibalić et al. (2017) This range of wavelengths offers exceptional precision in spectroscopic measurements, Prozument et al. (2011); Gordon et al. (2014) showing transient amplifiers of microwave radiation Raimond et al. (1982); Moi et al. (1983); Goy et al. (1983) and contributing to the development of the fifth generation mobile communication network.Deb and Kjærgaard (2018); Yuan et al. (2023b) It serves a crucial role in the advancement of atomic physics and quantum science research.Carter and Martin (2013); Kiffner et al. (2016); Grimes et al. (2017); Wade et al. (2017); Peper et al. (2019); Suleymanzade et al. (2020); Kanungo et al. (2022); Legaie, Raithel, and Anderson (2023)
Figure 1: (a) The numerical data demonstrates the scaling laws for the average orbit radius of Rydberg atoms in the $S$ orbit (blue circles) and circular orbit (red squares), expressed in terms of the effective principal quantum number denoted as $n^{*}$ for each case. (b) The scaling laws govern the radial dipole matrix elements for transitions between neighboring Rydberg states, denoted as $|\langle nP_{3/2}|e\hat{r}|nS_{1/2}\rangle|$ (blue diamonds), as well as for transitions between a low-lying state and a Rydberg state, represented as $|\langle nP_{3/2}|e\hat{r}|5S_{1/2}\rangle|$ (red hexagons), where $n^{*}=n-\delta_{n,2,j}$ for the $P$ orbit.
The expectation value of the electron orbit, determined by the radial wave function for an alkali atom state with medium to high quantum numbers $n$, is given byLindgård and Nielsen (1977); Theodosiou (1984)
@@
where $a_{0}\approx 5.29\times 10^{-2}~{}\rm~{}nm$ represents the Bohr radius. This formula offers an exact description of the radius of Rydberg atomic orbitals with high principal quantum numbers, as the Coulomb potential predominates in this regime. The Rydberg atomic wave function is influenced by quantum defects. They mainly generate a phase shift to the wave function. Hence wave functions of alkali atoms are different from that of hydrogen atoms as long as quantum defects are non-negligible. As a result, the expectation value of physical quantities can be directly derived from the hydrogen wave function, with the only alteration being the substitution of the principal quantum number $n$ with the effective principal quantum number $n^{*}$.Kazimierczuk et al. (2014); Heckötter et al. (2017) For low $l$ states, particularly $l=0$ ($S$ states), the expectation value simplifies to be $23n^{*2}a_{0}/$ 2, but for the circular orbit with $l=n-1$, where the quantum defect is zero, the value becomes $(n^{2}+n/2)a_{0}$. For large $n$ the radius of the Rydberg electron can be hundreds of nanometers. This permits the formation of ultra-long range Rydberg molecules, Herzberg (1987); Greene, Dickinson, and Sadeghpour (2000); Bendkowsky et al. (2009, 2010); Kiffner et al. (2012); Krupp et al. (2014); Camargo et al. (2016); Li et al. (2011); Eiles and Greene (2017); Eiles et al. (2017); Shaffer, Rittenhouse, and Sadeghpour (2018); Whalen et al. (2020); Giannakeas et al. (2020); Peper and Deiglmayr (2021); Christian Fey and Schmelcher (2020); Hummel, Eiles, and Schmelcher (2021) where ground state atoms are trapped within the Rydberg electron cloud. In Fig. 1(a), the ARC toolbox Šibalić et al. (2017) is used to perform numerical calculation of the expectation values for the $S$ and circular orbits of Rubidium, encompassing the principal quantum number $n$ from 10 to 100. The data exhibit quadratic scaling with $n^{*}$, closely aligning with the analytical expression.
@@
Lifetimes of Rydberg atoms are much longer than typical time scales of laser coupling and Rydberg atom interactions. This is important for maintaining quantum coherence, a crucial factor for quantum computation and simulation. The spontaneous decay rate from $|n,l,j\rangle$ to lower energy state $|n^{\prime},l^{\prime},j^{\prime}\rangle$ is governed by the Einstein coefficient (in SI units):
@@
where $\omega_{n^{\prime}l^{\prime}j^{\prime},nlj}$ denotes the corresponding transition angular frequency, $l_{max}$ signifies the larger of the two values between $l$ and $l^{\prime}$, and the curly braces indicate a Wigner-$6j$ symbol.Seiler et al. (2016); Beterov et al. (2009) The radial part of Rydberg electronic wave functions has a vanishing overlap with that of ground and weakly excited states, decreasing $A_{nlj\rightarrow n^{\prime}l^{\prime}j^{\prime}}$. At zero temperature, the overall decay rate of Rydberg states to many low-lying states is proportional to $n^{*-3}$, where in this case $\omega_{n^{\prime}l^{\prime}j^{\prime},nlj}$ is nearly constant and only the electric dipole transitions dominate. However, in circular Rydberg states ($l=n-1$), because $A_{nlj\rightarrow n^{\prime}l^{\prime}j^{\prime}}\propto\omega^{3}_{n^{\prime}l^{\prime}j^{\prime},nlj}$, they can only undergo a transition to an adjacent Rydberg state via a dipole-allowed process.Hulet and Kleppner (1983); Nguyen et al. (2018); Hölzl et al. (2024) This leads to the spontaneous emission rate of the circular state being proportional to $n^{*-5}$.
@@
Compared to the high $l$ states, the quantum defects of the low $l\leq 3$ states have a greater impact on the energy levels correction. Consequently, the weak external static electric field will not cause the Rydberg atom in the low orbit state to endure a direct transition between the adjacent orbits but will instead present the energy level shift as a second-order perturbation, which is in the form of a quadratic Stark effectKhadjavi, Lurio, and Happer (1968); Angel and Sandars (1968)
@@
$\Delta E_{\rm Stark}=\sum_{n^{\prime},l^{\prime}j^{\prime}\neq n,l,j}\frac{|\langle n^{\prime},l^{\prime},j^{\prime},m_{j}^{\prime}|e\hat{r}\cdot\bm{E}|n,l,j,m_{j}\rangle|^{2}}{E_{n^{\prime},l^{\prime},j^{\prime}}-E_{n,l,j}}=-\frac{1}{2}\alpha_{0}E^{2}$ (9)
@@
where we have assumed the static electric field $\bm{E}$ is applied along the $z$-axis creating off-diagonal couplings between states, with the selection rule $\Delta l=\pm 1$, $\Delta m_{j}=0$. As determined by the Wigner-Eckart theorem,Weissbluth (1978) the dipole matrix elements are linked to the radial matrix elements $\langle n^{\prime},l^{\prime},j^{\prime}|e\hat{r}|{n},l,j\rangle$ through coefficients independent of the principal quantum number. Therefore, the scalar polarizability $\alpha_{0}$ scales with $n^{*7}$ due to the combined scaling of the radial dipole matrix element and the energy spacing for neighboring Rydberg states in the denominator.Zimmerman et al. (1979); O’Sullivan and Stoicheff (1985); Lai et al. (2018); Bai et al. (2020a); Chakraborty and Sahoo (2023) It should be noted that states possessing quantum numbers of angular momentum $l>3$ manifest degeneracy, leading to a linear Stark shift. For a more precise description of the Stark structure of the Rydberg states in alkali atoms, direct diagonalization of the energy matrix is required. The Rydberg atoms due to high polarizability can be exploited as highly accurate sensors for electric fieldsSedlacek et al. (2012); Fan et al. (2014, 2015); Anderson, Paradis, and Raithel (2018); Cox et al. (2018); Meyer et al. (2020); Jing et al. (2020); Crawford et al. (2021); Meyer et al. (2021); Calero et al. (2021); Hu et al. (2022); Liu et al. (2022); Fancher et al. (2023); Liu et al. (2023a); Ding et al. (2022); Yuan et al. (2023a) and the nondestructive imaging of atoms and ions.Günter et al. (2012); Gross, Vogt, and Li (2020)
@@
Table 2: Scaling of various properties of Rydberg atoms with the effective principal quantum numbers $n^{*}$. The data for the transition rate in the parentheses is scaling in the circular state.